# Optimizing Posterior Samples for Bayesian Optimization via Rootfinding

### v1

@@ Abstract
Bayesian optimization devolves the global optimization of a costly objective function to the global optimization of a sequence of acquisition functions. This inner-loop optimization can be catastrophically difficult if it involves posterior samples, especially in higher dimensions. We introduce an efficient global optimization strategy for posterior samples based on global rootfinding. It provides gradient-based optimizers with judiciously selected starting points, designed to combine exploitation and exploration. The algorithm scales practically linearly to high dimensions. For posterior sample-based acquisition functions such as Gaussian process Thompson sampling (GP-TS) and variants of entropy search, we demonstrate remarkable improvement in both inner- and outer-loop optimization, surprisingly outperforming alternatives like EI and GP-UCB in most cases. We also propose a sample-average formulation of GP-TS, which has a parameter to explicitly control exploitation and can be computed at the cost of one posterior sample. Our implementation is available at https://github.com/UQUH/TSRoots.
@@ 1 Introduction
Some acquisition functions involve sample functions from the posterior model, which need to be optimized globally. Gaussian process Thompson sampling (GP-TS) [8] uses posterior samples directly as random acquisition functions. In many information-theoretic acquisition functions such as predictive entropy search (PES) [11], max-value entropy search (MES) [12], and joint entropy search (JES) [13, 14], multiple posterior samples are drawn and optimized to find their global optimum location and/or value. Such acquisition functions are celebrated for their nice properties in BO: TS has strong theoretical guarantees [7, 15] and can be scaled to high dimensions [16]; information-theoretic acquisition functions are grounded in principles for optimal experimental design [17]; and both types can be easily parallelized in synchronous batches [18, 19, 20]. However, posterior samples are much more complex than other acquisition functions as they fluctuate throughout the design space, and are less smooth than the posterior mean and marginal variance. The latter are the basis of many acquisition functions, such as expected improvement (EI) [1], probability of improvement (PI) [21], and upper confidence bound (GP-UCB) [5]. As a consequence, posterior samples have many more local optima, and the number scales exponentially with the input dimension.
@@
  • A novel strategy for the global optimization of posterior sample-based acquisition functions. The starting points are dependent on the posterior sample, so that each is close to a local optimum that is a candidate for the global optimum. The selection algorithm scales linearly to high dimensions.
@@ 2 General Background
Given a data set $\mathcal{D}=\{(\mathbf{x}^{i},y^{i})\}_{i=1}^{n}$, the posterior model $f|\mathcal{D}$ is also a GP. Samples from the posterior have a decoupled representation called pathwise conditioning, originally proposed in [30]:
@@ 3 Global Optimization of Posterior Samples
In this section, we introduce an efficient algorithm for the global optimization of posterior sample-based acquisition functions. For this, we exploit the separability of prior samples and useful properties of posterior samples to judiciously select a set of starting points for gradient-based multistart optimizers.
@@ 3.1 TS-roots Algorithm
To find the global minimum $(\widetilde{\mathbf{x}}_{\widetilde{\omega}}^{\star},\widetilde{f}_{\widetilde{\omega}}^{\star})$ of a posterior sample $\widetilde{f}_{\widetilde{\omega}}(\mathbf{x})$, we observe that given the assumptions, Eq. 1 can be rewritten as:
@@
Here, the prior sample $f(\mathbf{x})$ is a separable function determined by the random bits $\omega$. Data adjustment $b(\mathbf{x})$ is a sum of the canonical basis with coefficients $\mathbf{v}=(\mathbf{K}_{n,n}+\boldsymbol{\Sigma})^{-1}(\mathbf{y}-\mathbf{f}_{n}-\boldsymbol{\varepsilon})$. Both the data adjustment and the posterior sample are determined by the random bits $\widetilde{\omega}=(\omega,\boldsymbol{\varepsilon})$. In the following, we denote the prior and the posterior samples as $f_{\omega}$ and $\widetilde{f}_{\widetilde{\omega}}$, respectively.
@@
We define TS-roots as a global optimization algorithm for GP posterior samples (given the assumptions) via gradient-based multistart optimization, where the starting points include a subset of the local minima $\breve{X}_{\omega}$ of a corresponding prior sample, $S_{\text{e}}\subseteq\breve{X}_{\omega}$, and a subset of the observed locations, $S_{\text{x}}\subseteq X$. We call $S_{\text{e}}$ the exploration set and $S_{\text{x}}$ the exploitation set. Let
@@
be the $n_{\text{o}}$ smallest local minima of the prior sample and $\widetilde{\mu}(\mathbf{x})=\boldsymbol{\kappa}_{\cdot,n}(\mathbf{x})(\mathbf{K}_{n,n}+\boldsymbol{\Sigma})^{-1}\mathbf{y}$ be the posterior mean function. The starting points are defined as:
$S=S_{\text{e}}\cup S_{\text{x}},\quad S_{\text{e}}=\operatorname*{arg\,mink}_{\mathbf{x}\in S_{\text{o}}}(\widetilde{f}_{\widetilde{\omega}}(\mathbf{x}),n_{\text{e}}),\quad S_{\text{x}}=\operatorname*{arg\,mink}_{\mathbf{x}\in X}(\widetilde{\mu}(\mathbf{x}),n_{\text{x}}).$ (4)
Here, $n_{\text{e}}$ and $n_{\text{x}}$ are imposed to control the cost of gradient-based multistart optimization, and $n_{\text{o}}$ is set to limit the number of evaluations of $\widetilde{f}_{\widetilde{\omega}}$ in the determination of $S_{\text{e}}$. Considering the cost difference, in general we can have $n_{\text{o}}\gg n_{\text{e}}$.
@@ 3.4 Ordering of Prior Sample Local Minima
While the size of $\breve{X}_{\omega}$ grows exponentially in domain dimension $d$, its representation in Eq. 5 enables an efficient algorithm to compute the best subset $S_{\text{o}}$ (eq. 3) without enumerating its elements. With Eq. 5, we see that $\breve{X}_{\omega}^{-}$ consists of all the local minima of $f_{\omega}$ with negative function values. Consider the case where $\breve{X}_{\omega}^{-}$ has at least $n_{\text{o}}$ elements so that in the definition of $S_{\text{o}}$ we can replace $\breve{X}_{\omega}$ with $\breve{X}_{\omega}^{-}$, which in turn can be replaced with $\Xi^{(1)}$. As we will show later, the problem of finding the largest elements of $|f_{\omega}(\mathbf{x})|$ within $\Xi^{(1)}$ is easier than finding the smallest negative elements of $f_{\omega}(\mathbf{x})$. Once the former is solved, we can solve the latter simply by removing the positive elements. Therefore, we convert the problem of Eq. 3 into three steps:
@@ 4 Related Works
While their global optima guarantee the Bayes’ decision rule, BO acquisition functions are highly non-convex and difficult to optimize [9]. Nevertheless, less attention has been given to the development of robust algorithms for optimizing these acquisition functions. For this inner-loop optimization, gradient-based optimizers are often selected because of their fast convergence and robust performance [38]. The implementation of such optimizers is facilitated by Monte Carlo (MC) acquisition functions whose derivatives are easy to evaluate [9]. Gradient-based optimizers also allow multistart settings that use a set of starting points which can be, for example, midpoints of data points [39], uniformly distributed samples over the input variable space [3, 40], or random points from a Latin hypercube design [41]. However, multistart-based methods with random search may have difficulty determining the non-flat regions of acquisition functions, especially in high dimensions [10].
@@
As discussed in Section 1, the family of posterior sample-based acquisition functions is determined from samples of the posterior. GP-TS [8] is a notable member that extends the classical TS for finite-armed bandits to continuous settings of BO (see algorithms in Appendix D). GP-TS prefers exploration by the mechanism that iteratively samples a function from the GP posterior of the objective function, optimizes this function, and selects the resulting solution as the next candidate for objective evaluation. To further improve the exploitation of GP-TS, the sample mean of MC acquisition functions can be defined from multiple samples of the posterior [9, 27]. Different types of MC acquisition functions can also be used to inject beliefs about functions into the prior [42].
@@ 5 Results
We assess the performance of TS-roots in optimizing benchmark functions. We then compare the quality of solutions to the inner-loop optimization of GP-TS acquisition functions obtained from our proposed method, a gradient-based multistart optimizer with uniformly random starting points, and a genetic algorithm. We also show how TS-roots can improve the performance of MES. Finally, we propose a new sample-average posterior function and show how it affects the performance of GP-TS. The experimental details for the presented results are in Appendix F.
@@
We test the empirical performance of TS-roots on challenging minimization problems of five benchmark functions: the 2D Schwefel, 4D Rosenbrock, 10D Levy, 16D Ackley, and 16D Powell functions [43]. The analytical expressions for these functions and their global minimum are given in Appendix E.
@@
Figure 2 shows the medians and interquartile ranges of solution values obtained from 20 runs of each of the considered BO methods. The corresponding histories of solution locations are in Figure 7 of Appendix G. With a fair comparison of outer-loop results (detailed in Appendix F), TS-roots surprisingly shows the best performance on the 2D Schwefel, 16D Ackley, and 16D Powell functions, and gives competitive results in the 4D Rosenbrock and 10D Levy problems. Notably, TS-roots recommends better solutions than its counterparts, TS-DSRF and TS-RF, in high-dimensional problems and offers competitive performance in low-dimensional problems. Across all the examples, EI and LCB tend to perform better in the initial stages, while TS-roots shows fast improvement in later stages. This is because GP-TS favors exploration, which delays rewards. The exploration phase, in general, takes longer for higher-dimensional problems.
@@
We assess the quality of solutions and computational cost for the inner-loop optimization of GP-TS acquisition functions by the proposed global optimization algorithm, termed rootfinding hereafter. We do so by computing the optimized values $\alpha_{k}^{\star}$ of the GP-TS acquisition functions, the corresponding solution points $\mathbf{x}_{k}^{\star}$, and the CPU times $t_{k}$ required for optimizing the acquisition functions during the optimization process. For low-dimensional problems of the 2D Schwefel and 4D Rosenbrock functions, we also compute the exact global solution points $\mathbf{x}_{k}^{\mathrm{t}}$ of the GP-TS acquisition functions by starting the gradient-based optimizer at a large number of initial points (set as $10^{4}$), which is much greater than the maximum possible number of starting points set for TS-roots. For comparison, we extend the same GP-TS acquisition functions to inner-loop optimization using a gradient-based multistart optimizer with random starting points (i.e., random multistart) and a genetic algorithm. In each outer-loop optimization iteration, the number of starting points for the random multistart and the population size of the genetic algorithm are equal to the number of starting points recommended for rootfinding. The same termination conditions are used for the three algorithms.
Figure 3 shows the comparative performance of the inner-loop optimization for low-dimensional cases: the 2D Schwefel and 4D Rosenbrock functions. We see that the optimized acquisition function values and the optimization runtimes by rootfinding and the random multistart algorithm are almost identical, both of which are much better than those by the genetic algorithm. Rootfinding gives the best quality of the new solution points in both cases, while the genetic algorithm gives the worst. Notably in higher-dimensional settings of the 10D Levy, 16D Ackley, and 16D Powell functions shown in Figure 4, rootfinding performs much better than the random multistart and genetic algorithm in terms of optimized acquisition values and optimization runtimes, which verifies the importance of the judicious selection of starting points for global optimization of the GP-TS acquisition functions and the efficiency of rootfinding in high dimensions. The random multistart becomes worse in high dimensions.
@@
We show how TS-roots can enhance the performance of MES [12], which uses information about the maximum function value $f^{\star}$ for processing BO. One approach to computing MES generates a set of GP posterior samples using TS-RF and subsequently optimizes the generated functions for samples of $f^{\star}$ using a gradient-based multistart optimizer with a large number of random starting points [12]. We hypothesize that high-quality $f^{\star}$ samples can improve the performance of MES. Thus, we assign both TS-roots and TS-RF as the inner workings of MES and then compare the resulting optimal solutions.
@@
Figure 5 shows the optimization histories for ten independent runs of each MES method. On the 4D Rosenbrock and 6d Hartmann functions, MES with TS-roots demonstrates superior optimization performance and faster convergence compared to MES with TS-RF, especially when 50 samples of $f^{\star}$ are generated. For the 10D Levy function, TS-roots outperforms TS-RF when using 10 samples of $f^{\star}$, while their performances are comparable when 50 samples are used.
Sample-Average Posterior Function.
[…]
We investigate how $\alpha_{\text{aTS}}(\mathbf{x})$ influences the outer-loop optimization results. For this, we set $N_{\text{c}}\in\{1,10,50,100\}$ for TS-roots to optimize the 2D Schwefel, 4D Rosenbrock, and 6d Ackley functions. We observe that increasing $N_{\text{c}}$ from $1$ to $10$ improves TS-roots performance on the 2D Schwefel, 4D Rosenbrock, and 6d Ackley functions (see Figure 9 in Appendix G). However, further increases in $N_{\text{c}}$ from 10 to 50 and 100 result in slight declines in solution quality as TS-roots transitions to exploitation. These observations indicate that there is an optimal value of $N_{\text{c}}$ for each problem at which TS-roots achieves its best performance by balancing exploitation and exploration. However, identifying such an optimal value to maximize the performance of $\alpha_{\text{aTS}}(\mathbf{x})$ for a particular optimization problem is an open issue.

### v2

@@ Abstract
Bayesian optimization devolves the global optimization of a costly objective function to the global optimization of a sequence of acquisition functions. This inner-loop optimization can be catastrophically difficult if it involves posterior sample paths, especially in higher dimensions. We introduce an efficient global optimization strategy for posterior samples based on global rootfinding. It provides gradient-based optimizers with two sets of judiciously selected starting points, designed to combine exploration and exploitation. The number of starting points can be kept small without sacrificing optimization quality. Remarkably, even with just one point from each set, the global optimum is discovered most of the time. The algorithm scales practically linearly to high dimensions, breaking the curse of dimensionality. For Gaussian process Thompson sampling (GP-TS), we demonstrate remarkable improvement in both inner- and outer-loop optimization, surprisingly outperforming alternatives like EI and GP-UCB in most cases. Our approach also improves the performance of other posterior sample-based acquisition functions, such as variants of entropy search. Furthermore, we propose a sample-average formulation of GP-TS, which has a parameter to explicitly control exploitation and can be computed at the cost of one posterior sample. Our implementation is available at https://github.com/UQUH/TSRoots.
@@ 1 Introduction
Some acquisition functions involve sample functions from the posterior model, which need to be optimized globally. Gaussian process Thompson sampling (GP-TS) [8] uses posterior sample paths directly as random acquisition functions. In many information-theoretic acquisition functions such as entropy search (ES) [11], predictive entropy search (PES) [12], max-value entropy search (MES) [13], and joint entropy search (JES) [14, 15], multiple posterior samples are drawn and optimized to find their global optimum location and/or value. Such acquisition functions are celebrated for their nice properties in BO: TS has strong theoretical guarantees [7, 16] and can be scaled to high dimensions [17]; information-theoretic acquisition functions are grounded in principles for optimal experimental design [18]; and both types can be easily parallelized in synchronous batches [19, 20, 21]. However, posterior sample paths are much more complex than other acquisition functions, as they fluctuate throughout the design space, and are less smooth than the posterior mean and marginal variance. The latter are the basis of many acquisition functions, such as expected improvement (EI) [1], probability of improvement (PI) [22], and upper confidence bound (GP-UCB) [5]. As a consequence, posterior sample paths have many more local optima, and the number scales exponentially with the input dimension.
@@
  • A novel strategy for the global optimization of posterior sample paths. The starting points are dependent on the posterior sample, so that each is close to a local optimum that is a candidate for the global optimum. The selection algorithm scales linearly to high dimensions.
@@ 2 General Background
Given a dataset $\mathcal{D}=\{(\mathbf{x}^{i},y^{i})\}_{i=1}^{n}$, the posterior model $f|\mathcal{D}$ is also a GP. Samples from the posterior have a decoupled representation called pathwise conditioning, originally proposed in [31, 32]:
@@ 3 Global Optimization of Posterior Samples
In this section, we introduce an efficient algorithm for the global optimization of posterior sample paths. For this, we exploit the separability of prior samples and useful properties of posterior samples to judiciously select a set of starting points for gradient-based multistart optimizers.
@@
Algorithm 1 TS-roots: Optimizing posterior samples via rootfinding
1:prior samples $f(\mathbf{x})$, $\boldsymbol{\varepsilon}$; prior covariances $\kappa(\mathbf{x},\mathbf{x}^{\prime})$, $\boldsymbol{\Sigma}$; dataset $\mathcal{D}$; set sizes $n_{\text{o}},n_{\text{e}},n_{\text{x}}$.
2:$S_{\text{o}}\leftarrow\texttt{minsort}(f(\mathbf{x}),n_{\text{o}})$triangleright Smallest $n_{\text{o}}$ local minima of the prior sample (Algorithm 4)
3:$[\widetilde{\mathbf{f}}_{\text{e}},I_{\text{e}}]\leftarrow\texttt{mink}(\widetilde{f}(S_{\text{o}}),n_{\text{e}})$triangleright Smallest $n_{\text{e}}$ values and indices of the posterior sample in $S_{\text{o}}$
4:$[\widetilde{\mathbf{f}}_{\text{x}},I_{\text{x}}]\leftarrow\texttt{mink}(\widetilde{f}(X),n_{\text{x}})$triangleright Smallest $n_{\text{x}}$ values and indices of the posterior mean in $X$
5:$S_{\text{e}}\leftarrow S_{\text{o}}[I_{\text{e}},:],\;S_{\text{x}}\leftarrow X[I_{\text{x}},:]$triangleright Starting points: exploration set and exploitation set
6:$[\widetilde{\mathbf{x}}^{\star},\widetilde{f}^{\star}]\leftarrow\texttt{minimize}(\widetilde{f}(\mathbf{x}),S),\;S=S_{\text{e}}\cup S_{\text{x}}$triangleright Gradient-based multistart optimization
7:Thompson sample $\widetilde{\mathbf{x}}^{\star}$triangleright Global minimum of the posterior sample
@@ 3.1 TS-roots Algorithm
We observe that, given the assumptions, a posterior sample in Eq. 1 can be rewritten as:
@@
Here, the prior sample $f(\mathbf{x})$ is a separable function determined by the random bits $\omega$. Data adjustment $b(\mathbf{x})$ is a sum of the canonical basis with coefficients $\mathbf{v}=(\mathbf{K}_{n,n}+\boldsymbol{\Sigma})^{-1}(\mathbf{y}-\mathbf{f}_{n}-\boldsymbol{\varepsilon})$. Both the data adjustment and the posterior sample are determined by the random bits $\widetilde{\omega}=(\omega,\boldsymbol{\varepsilon})$. In the following, we denote the prior and the posterior samples as $f_{\omega}$ and $\widetilde{f}_{\widetilde{\omega}}$, respectively. Our goal is to find the global minimum $(\widetilde{\mathbf{x}}_{\widetilde{\omega}}^{\star},\widetilde{f}_{\widetilde{\omega}}^{\star})$ of the posterior sample $\widetilde{f}_{\widetilde{\omega}}(\mathbf{x})$.
@@
We define TS-roots as a global optimization algorithm for GP posterior samples (given the assumptions) via gradient-based multistart optimization. The starting points include: (1) a subset $S_{\text{e}}$ of the local minima $\breve{X}_{\omega}$ of a corresponding prior sample; and (2) a subset $S_{\text{x}}$ of the observed locations $X$. We call $S_{\text{e}}$ the exploration set and $S_{\text{x}}$ the exploitation set. Specifically, let
@@
be the $n_{\text{o}}$ smallest local minima of the prior sample. The set of starting points, $S$, is defined as:
$S=S_{\text{e}}\cup S_{\text{x}},\quad S_{\text{e}}=\operatorname*{arg\,mink}_{\mathbf{x}\in S_{\text{o}}}(\widetilde{f}_{\widetilde{\omega}}(\mathbf{x}),n_{\text{e}}),\quad S_{\text{x}}=\operatorname*{arg\,mink}_{\mathbf{x}\in X}(\widetilde{f}_{\widetilde{\omega}}(\mathbf{x}),n_{\text{x}}).$ (4)
Algorithm 1 outlines the procedure for TS-roots. Here, $n_{\text{e}}$ and $n_{\text{x}}$ are imposed to control the cost of gradient-based multistart optimization, and $n_{\text{o}}$ is set to limit the number of evaluations of $\widetilde{f}_{\widetilde{\omega}}$ in the determination of $S_{\text{e}}$. Considering the cost difference, we can have $n_{\text{o}}\gg n_{\text{e}}$. We observe that $n_{\text{e}}$ and $n_{\text{x}}$ can be set to small values, and $n_{\text{o}}$ to a moderate value, without sacrificing the quality of optimization, see Appendix D. The TS-roots algorithm scales linearly in $d$, see Section C.5.
@@ 3.4 Ordering of Prior Sample Local Minima
While the size of $\breve{X}_{\omega}$ grows exponentially in domain dimension $d$, its representation in Eq. 5 enables an efficient algorithm to compute the best subset $S_{\text{o}}$ (Eq. 3) without enumerating its elements.
With Eq. 5, we see that $\breve{X}_{\omega}^{-}$ consists of all the local minima of $f_{\omega}$ with negative function values. Consider the case where $\breve{X}_{\omega}^{-}$ has at least $n_{\text{o}}$ elements so that in the definition of $S_{\text{o}}$ we can replace $\breve{X}_{\omega}$ with $\breve{X}_{\omega}^{-}$, which in turn can be replaced with $\Xi^{(1)}$. As we will show later, the problem of finding the largest elements of $|f_{\omega}(\mathbf{x})|$ within $\Xi^{(1)}$ is easier than finding the smallest negative elements of $f_{\omega}(\mathbf{x})$. Once the former is solved, we can solve the latter simply by removing the positive elements. Therefore, we convert the problem of Eq. 3 into three steps:
@@ 5 Related Works
While their global optima guarantee the Bayes’ decision rule, BO acquisition functions are highly non-convex and difficult to optimize [9]. Nevertheless, less attention has been given to the development of robust algorithms for optimizing these acquisition functions. For this inner-loop optimization, gradient-based optimizers are often selected because of their fast convergence and robust performance [41]. The implementation of such optimizers is facilitated by Monte Carlo (MC) acquisition functions whose derivatives are easy to evaluate [9]. Gradient-based optimizers also allow multistart settings that use a set of starting points which can be, for example, midpoints of data points [42], uniformly distributed samples over the input variable space [3, 43], or random points from a Latin hypercube design [44]. However, multistart-based methods with random search may have difficulty determining the non-flat regions of acquisition functions, especially in high dimensions [10]. The log reformulation approach is a good solution to the numerical pathology of flat acquisition surface over large regions of the input variable space [43]. While this approach works for acquisition functions prone to the flat surface issue such as the family of EI-based acquisition functions, its performance has yet to be evaluated for acquisition functions with many local minima like those based on posterior samples.
@@
As discussed in Section 1, the family of posterior sample-based acquisition functions is determined from samples of the posterior. GP-TS [8] is a notable member that extends the classical TS for finite-armed bandits to continuous settings of BO (see algorithms in Appendix E). GP-TS prefers exploration by the mechanism that iteratively samples a function from the GP posterior of the objective function, optimizes this function, and selects the resulting solution as the next candidate for objective evaluation. To further improve the exploitation of GP-TS, the sample mean of MC acquisition functions can be defined from multiple samples of the posterior [9, 28]. Different types of MC acquisition functions can also be used to inject beliefs about functions into the prior [45].
@@ 6 Results
We assess the performance of TS-roots in optimizing benchmark functions. We then compare the quality of solutions to the inner-loop optimization of GP-TS acquisition functions obtained from our proposed method, a gradient-based multistart optimizer with uniformly random starting points, and a genetic algorithm. We also show how TS-roots can improve the performance of MES. Finally, we propose a new sample-average posterior function and show how it affects the performance of GP-TS. The experimental details for the presented results are in Appendix H.
@@
We test the empirical performance of TS-roots on challenging minimization problems of five benchmark functions: the 2D Schwefel, 4D Rosenbrock, 10D Levy, 16D Ackley, and 16D Powell functions [46]. The analytical expressions for these functions and their global minimum are given in Appendix F.
@@
Figure 2 shows the medians and interquartile ranges of solution values obtained from 20 runs of each of the considered BO methods. The corresponding histories of solution locations are in Figure 9 of Appendix I. With a fair comparison of outer-loop results (detailed in Appendix H), TS-roots surprisingly shows the best performance on the 2D Schwefel, 16D Ackley, and 16D Powell functions, and gives competitive results in the 4D Rosenbrock and 10D Levy problems. Notably, TS-roots recommends better solutions than its counterparts, TS-DSRF and TS-RF, in high-dimensional problems and offers competitive performance in low-dimensional problems. Across all the examples, EI and LCB tend to perform better in the initial stages, while TS-roots shows fast improvement in later stages. This is because GP-TS favors exploration, which delays rewards. The exploration phase, in general, takes longer for higher-dimensional problems.
@@
Optimizing Real-world Problem.
We implement TS-roots to optimize an engineered ten-bar truss structure (see Appendix G). Ten design variables of the truss are the cross-sectional areas of the truss members. The objective is to minimize a weighted sum of the scaled total cross-sectional area and the scaled vertical displacement at a node of interest.
Figure 2(f) shows the outer-loop optimization results for the truss obtained from 20 runs of each BO method, where $f^{\star}$ is a lower bound of the best objective function value we observed from all runs. TS-roots provides the best optimization result with rapid convergence.
@@
We assess the quality of solutions and computational cost for the inner-loop optimization of GP-TS acquisition functions by the proposed global optimization algorithm, referred to as rootfinding hereafter. We do so by computing the optimized values $\alpha_{k}^{\star}$ of the GP-TS acquisition functions, the corresponding solution points $\mathbf{x}_{k}^{\star}$, and the CPU times $t_{k}$ required for optimizing the acquisition functions during the optimization process. For low-dimensional problems of the 2D Schwefel and 4D Rosenbrock functions, we also compute the exact global solution points $\mathbf{x}_{k}^{\mathrm{t}}$ of the GP-TS acquisition functions by starting the gradient-based optimizer at a large number of initial points (set as $10^{4}$), which is much larger than the maximum number of starting points set for TS-roots. For comparison, we extend the same GP-TS acquisition functions to inner-loop optimization using a gradient-based multistart optimizer with random starting points (i.e., random multistart) and a genetic algorithm. In each outer-loop optimization iteration, the number of starting points for the random multistart and the population size of the genetic algorithm are equal to the number of starting points recommended for rootfinding. The same termination conditions are used for the three algorithms.
Figure 3 shows the comparative performance of the inner-loop optimization for low-dimensional cases: the 2D Schwefel and 4D Rosenbrock functions. We see that the optimized acquisition function values and the optimization runtimes by rootfinding and the random multistart algorithm are almost identical, both of which are much better than those by the genetic algorithm. Rootfinding gives the best quality of the new solution points in both cases, while the genetic algorithm gives the worst. Notably in higher-dimensional settings of the 10D Levy, 16D Ackley, and 16D Powell functions shown in Figure 4, rootfinding performs much better than the random multistart and genetic algorithm in terms of optimized acquisition values and optimization runtimes, which verifies the importance of the judicious selection of starting points for global optimization of the GP-TS acquisition functions and the efficiency of rootfinding in high dimensions. The performance of random multistart becomes worse in higher dimensions. Appendix I provides additional results for gradient-based multistart optimization using two other initialization schemes: uniform grid and Latin hypercube sampling. Rootfinding outperforms both, especially in higher dimensions.
@@
We show how TS-roots can enhance the performance of MES [13], which uses information about the maximum function value $f^{\star}$ for conducting BO. One approach to computing MES generates a set of GP posterior samples using TS-RF and subsequently optimizes the generated functions for samples of $f^{\star}$ using a gradient-based multistart optimizer with a large number of random starting points [13]. We hypothesize that high-quality $f^{\star}$ samples can improve the performance of MES. Thus, we assign both TS-roots and TS-RF as the inner workings of MES and then compare the resulting optimal solutions. Note that the inner-loop optimization of MES, which strongly influences the optimization results, is not addressed by TS-roots.
@@
Figure 5 shows the optimization histories for ten independent runs of each MES method. On the 4D Rosenbrock and 6D Hartmann functions, MES with TS-roots demonstrates superior optimization performance and faster convergence compared to MES with TS-RF, especially when 50 samples of $f^{\star}$ are generated. For the 10D Levy function, TS-roots outperforms TS-RF when using 10 samples of $f^{\star}$, while their performance is comparable when 50 samples are used.
Performance of Sample-Average Posterior Functions.
We investigate how $\alpha_{\text{aTS}}(\mathbf{x})$ influences the outer-loop optimization results. For this, we set $N_{\text{c}}\in\{1,10,50,100\}$ for TS-roots to optimize the 2D Schwefel, 4D Rosenbrock, and 6D Ackley functions. We observe that increasing $N_{\text{c}}$ from $1$ to $10$ improves TS-roots performance on the 2D Schwefel, 4D Rosenbrock, and 6D Ackley functions (see Figure 13 in Appendix I). However, further increases in $N_{\text{c}}$ from 10 to 50 and 100 result in slight declines in solution quality as TS-roots transitions to exploitation. These observations indicate that there is an optimal value of $N_{\text{c}}$ for each problem at which TS-roots achieves its best performance by balancing exploitation and exploration. However, identifying the optimal value to maximize the performance of $\alpha_{\text{aTS}}(\mathbf{x})$ for a particular optimization problem remains an open issue.